Q3: A network is given as 190.28.0.0/16. Eight subnets are to be created.

(a) What will be the CIDR network prefix?

(b) Express last subnet in dotted decimal notation with mask.

(c) What is the maximum number of hosts in each subnet?

(d) What is the address range for hosts in second subnet?

(e) What is the subnet broadcast address for the first subnet?(10 marks)


1
Expert's answer
2022-03-24T16:21:31-0400

(a) "2^3=8" so 3 bits will be enough to create 8 subnets. The CIDR network prefix: /19.

(b) 190.28.224.0/19

(c) The maximum number of hosts in each subnet is "2^h-2", where h is a number of host bits.

"h=32-19=13"

"2^{13}-2=8190" hosts.

(d) 190.28.32.1 to 190.28.63.254 (it is clear if you write the address range in binary form)

(e) 190.28.31.255